# Recent Advances and Trends in Research Paper Recommender Systems: A Comprehensive Survey

### v1

@@ 4 Discussion
    $\circ$ On What are the relevance criteria used to evaluate the recommendations? (RQ3.1) Relevance criteria in RP-RSs have evolved over time. When comparing the data from the current survey with those of a previous survey [35], several notable differences emerge. Paper references are the most frequently used criterion in the current survey, accounting for 44.29% of the methods, compared to 24.07% in the previous survey. This increase likely reflects a growing reliance on citation-based relevance, driven by the increasing importance of citation metrics and the availability of citation-based datasets, which are more accessible and scalable compared to user interaction data. User interactions, including clicks, reads, and downloads, are utilized by 28.57% of the systems in the current survey, showing a slight decline from 31.48% in the earlier survey. This decrease can be attributed to the challenges of acquiring real user interaction data, particularly in large-scale settings, and the increasing concern around user privacy. Human judgment was employed in 8.57% of the methods in the current survey, compared to 27.78% in the previous survey. This decline reflects the resource-intensive nature of manual evaluation, leading to a preference for automated or semi-automated evaluation methods that can handle larger datasets more efficiently. Labeled papers, now at 12.86%, were previously included in the other category. Finally, the other category has seen a significant reduction, from 24.07% in the previous survey to just 4.29% in the current one. This shift indicates a movement toward more explicit, standardized, and structured evaluation approaches, as researchers are increasingly prioritizing well-defined, reproducible criteria. Overall, the increase in the use of paper references points to a growing focus on citation-based relevance, while the decline in user interactions and human judgment reflects the growing emphasis on scalability, data availability, and the challenges of manual evaluation.
@@
  • On What are the main challenges addressed by the RP-RSs? (RQ4). RP-RSs continue to face several key challenges that hinder their full potential. One prominent issue is the insufficient incorporation of user modeling, as many systems still rely on text-based content or specific papers rather than personalized user profiles, although there has been a 10% increase in the adoption of user modeling in recent years. Additionally, the overreliance on precision as the primary evaluation metric persists, despite growing recognition of the need for more comprehensive metrics such as recall, NDCG, novelty, and user satisfaction, which help ensure relevance, diversity, and temporal accuracy [8, 7, 36]. The gap between theoretical advancements and practical application remains significant, with only 9.52% of works focusing on real-world prototypes or systems, indicating scalability challenges. Similarly, the lack of detailed documentation persists, with many works failing to provide adequate transparency for replicability. Cold start and sparsity issues are ongoing, with a dominant focus on content-based methods to address these challenges, reflecting the continued complexity of these problems in real-world settings. Privacy concerns also remain underaddressed, as many works still overlook how to protect sensitive data and ensure proper anonymization. Serendipity, though crucial for fostering innovation, has not received sufficient attention, with few works explicitly focusing on it. The evaluation of RP-RSs remains limited to traditional accuracy metrics, necessitating a broader adoption of metrics that account for novelty, diversity, and user satisfaction, as highlighted by recent works in the field. Furthermore, the lack of a clearly defined target audience in many systems continues to hinder the design of user-centered recommendations, with only 12.69% of works explicitly considering audience differentiation. Lastly, dataset accessibility issues persist, with a significant portion of datasets being either proprietary, modified, or inaccessible, complicating reproducibility efforts. These ongoing challenges indicate that while progress has been made, there is still much to be addressed for RP-RSs to achieve their full potential.
@@ 5 Conclusion
In this survey, a deep and detailed analysis of RP-RSs from November 2021 to December 2024, building upon the previous surveys by [9, 6, 35] is provided. A total of 63 works have been analyzed in depth, providing a comprehensive snapshot of the current trends and challenges in RP-RSs. Unlike prior surveys, this survey goes beyond simply examining the techniques and models, offering a through exploration of how these methods are applied at various stages of the recommendation process, which was not fully addressed in previous surveys. Additionally, multiple tables that organize and summarize the key elements of RP-RSs are presented, providing valuable insights into the current status of the field. By addressing key research questions—such as how RP-RSs generate recommendations, the datasets used, evaluation methods, the challenges faced, and the emerging challenges in the field—a comprehensive understanding of the current status of RP-RSs is gained.

### v2

@@ 4 Discussion
    $\circ$ On What are the relevance criteria used to evaluate the recommendations? (RQ3.1) Relevance criteria in RP-RSs have evolved over time. When comparing the data from the current survey with those of a previous survey [IJ:DL:KREUTZ:2022], several notable differences emerge. Paper references are the most frequently used criterion in the current survey, accounting for 44.29% of the methods, compared to 24.07% in the previous survey. This increase likely reflects a growing reliance on citation-based relevance, driven by the increasing importance of citation metrics and the availability of citation-based datasets, which are more accessible and scalable compared to user interaction data. User interactions, including clicks, reads, and downloads, are utilized by 28.57% of the systems in the current survey, showing a slight decline from 31.48% in the earlier survey. This decrease can be attributed to the challenges of acquiring real user interaction data, particularly in large-scale settings, and the increasing concern around user privacy. Human judgment was employed in 8.57% of the methods in the current survey, compared to 27.78% in the previous survey. This decline reflects the resource-intensive nature of manual evaluation, leading to a preference for automated or semi-automated evaluation methods that can handle larger datasets more efficiently. Labeled papers, now at 12.86%, were previously included in the other category. Finally, the other category has seen a significant reduction, from 24.07% in the previous survey to just 4.29% in the current one. This shift indicates a movement toward more explicit, standardized, and structured evaluation approaches, as researchers are increasingly prioritizing well-defined, reproducible criteria. Overall, the increase in the use of paper references points to a growing focus on citation-based relevance, while the decline in user interactions and human judgment reflects the growing emphasis on scalability, data availability, and the challenges of manual evaluation.
@@
  • On What are the main challenges addressed by the RP-RSs? (RQ4). RP-RSs continue to face several key challenges that hinder their full potential. One prominent issue is the insufficient incorporation of user modeling, as many systems still rely on text-based content or specific papers rather than personalized user profiles, although there has been a 10% increase in the adoption of user modeling in recent years. Additionally, the overreliance on precision as the primary evaluation metric persists, despite growing recognition of the need for more comprehensive metrics such as recall, NDCG, novelty, and user satisfaction, which help ensure relevance, diversity, and temporal accuracy [IJ:DL:BEEL:2015, IJ:IEEE:BAI:2019, IJ:DL:KREUTZ:2022]. The gap between theoretical advancements and practical application remains significant, with only 9.52% of works focusing on real-world prototypes or systems, indicating scalability challenges. Similarly, the lack of detailed documentation persists, with many works failing to provide adequate transparency for replicability. Cold start and sparsity issues are ongoing, with a dominant focus on content-based methods to address these challenges, reflecting the continued complexity of these problems in real-world settings. Privacy concerns also remain underaddressed, as many works still overlook how to protect sensitive data and ensure proper anonymization. Serendipity, though crucial for fostering innovation, has not received sufficient attention, with few works explicitly focusing on it. The evaluation of RP-RSs remains limited to traditional accuracy metrics, necessitating a broader adoption of metrics that account for novelty, diversity, and user satisfaction, as highlighted by recent works in the field. Furthermore, the lack of a clearly defined target audience in many systems continues to hinder the design of user-centered recommendations, with only 12.69% of works explicitly considering audience differentiation. Lastly, dataset accessibility issues persist, with a significant portion of datasets being either proprietary, modified, or inaccessible, complicating reproducibility efforts. These ongoing challenges indicate that while progress has been made, there is still much to be addressed for RP-RSs to achieve their full potential.
@@ 5 Conclusion
In this survey, a deep and detailed analysis of RP-RSs from November 2021 to December 2024, building upon the previous surveys by [IJ:DL:BEEL:2015, IJ:IEEE:BAI:2019, IJ:DL:KREUTZ:2022] is provided. A total of 63 works have been analyzed in depth, providing a comprehensive snapshot of the current trends and challenges in RP-RSs. Unlike prior surveys, this survey goes beyond simply examining the techniques and models, offering a through exploration of how these methods are applied at various stages of the recommendation process, which was not fully addressed in previous surveys. Additionally, multiple tables that organize and summarize the key elements of RP-RSs are presented, providing valuable insights into the current status of the field. By addressing key research questions—such as how RP-RSs generate recommendations, the datasets used, evaluation methods, the challenges faced, and the emerging challenges in the field—a comprehensive understanding of the current status of RP-RSs is gained.